OpenStudy (anonymous):

A dress that costs $65.99 is on sale for 14 off. What is the sale price of the dress? Round your answer to the nearest cent. The sale price of the dress is $

OpenStudy (anonymous):

1/4 off*

OpenStudy (anonymous):

1/4? or 14?

OpenStudy (anonymous):

And is there a percent?

OpenStudy (anonymous):

If not, 1/4 of 65.99 is the same as dividing by 4. What's 65.99 / 4?

OpenStudy (anonymous):

16.49

OpenStudy (anonymous):

Yes, but if you are going to round it, we round up, not down. It should be 16.5 Now subtract 65.99 - 16.5

OpenStudy (anonymous):

49.49

OpenStudy (anonymous):

Yes, so that is the price.

OpenStudy (ahsome):

If it is one fourth, simply multiply by one-fourth \[Savings = 65.99 \times \frac{1}{4}\]\[Savings = \frac{65.99}{4}\]\[Savings =16.4975\] Now, take away the total price from the savings \[T.P. = Price - Savings\]\[T.P. = 65.99-16.4975\]\[T.P.=49.4925\]\[T.P.\approx$49.49\]